bill533:

Midpoint formula: M = x1 + x2 / 2, y1 + y2/2 What is the midpoint between point P and point Q? A) (2, 4) B) (3, 5) C) (4, 5) D) (4, 6) E) (5, 2)

bill533:

@Gdeinward

Gdeinward:

Post the points please. ;D

bill533:

1 attachment
Gdeinward:

so input the points and simplify? Have you attempted solving the problem?

bill533:

P = (-2,3) Q = (8,7) Add up the x coordinates -2+8 = 6 then divide by 2 6/2 = 3 Right away we can see that B is the answer as its the only choice that has 3 as the x coordinate. But lets keep going to finish off the problem


Add up the y coordinates 3+7 = 10 Divide by 2 10/2 = 5
Answer: Choice B) (3,5)